OpenStudy (anonymous):

The volume of a sphere is found using the formula V= 4/3 3.14r^3. if the radius of a sphere is 3/4 feet, what is the volume of the sphere in cubic feet?

OpenStudy (anonymous):

\[\frac{ 4 }{ 3 }\times3.14\times \left( \frac{ 3 }{ 4} \right)^3\]

OpenStudy (anonymous):

=0.9935